### v2

# Domain decomposition-based coupling of Operator Inference Reduced Order Models via the Schwarz alternating method

Abstract
This paper presents and evaluates an approach for coupling together subdomain-local reduced order models (ROMs) constructed via non-intrusive operator inference (OpInf) with each other and with subdomain-local full order models (FOMs), following a domain decomposition of the spatial geometry on which a given partial differential equation (PDE) is posed. Joining subdomain-local models is accomplished using the overlapping Schwarz alternating method, a minimally-intrusive multiscale coupling technique that works by transforming a monolithic problem into a sequence of subdomain-local problems, which communicate through transmission boundary conditions imposed on the subdomain interfaces. After formulating the overlapping Schwarz alternating method for OpInf ROMs, termed OpInf-Schwarz, we evaluate the method’s accuracy and efficiency on several test cases involving the heat equation in two spatial dimensions. We demonstrate that the method is capable of coupling together arbitrary combinations of OpInf ROMs and FOMs, and that speed-ups over a monolithic FOM are possible when performing OpInf ROM coupling.
1 Introduction
Despite advancements in both computer architectures and algorithms, the modeling and simulation of complex physical systems often requires tremendous computational resources. These requirements may preclude many-query analyses such as engineering design or uncertainty quantification. While projection-based reduced order models (ROMs) have promised to mitigate this difficulty, traditional intrusive ROMs, e.g., Galerkin [13, 37] and least squares Petrov-Galerkin (LSPG) projection ROMs [4], have their own shortcomings, including a lack of systematic refinement mechanisms, a lack of robustness, stability, and accuracy in the predictive regime, and lengthy implementation time requirements.
This paper presents a promising approach for mitigating the aforementioned difficulties by enabling domain decomposition- (DD-)based coupling of subdomain-local ROMs with each other and/or with subdomain-local full order models (FOMs). Our approach is based on the following ingredients: (i) a decomposition of the physical domain of interest into two or more overlapping subdomains, (ii) the construction of subdomain-local ROMs and/or FOMs in each of the subdomains, and (iii) the rigorous coupling of the subdomain-local models via the Schwarz alternating method [33]. The Schwarz alternating method is based on the simple idea that, if the solution to a partial differential equation (PDE) is known in two or more regularly shaped domains comprising a more complex domain, these local solutions can be used to iteratively build a solution for the more complex domain, with information propagating between the subdomains through carefully constructed transmission boundary conditions (BCs). We choose the Schwarz alternating method since it has a number of advantages over competing multiscale coupling methods. These advantages of the Schwarz alternating method include its concurrent nature, its ability to couple non-conformal meshes with different element topologies and different time integrators with different time steps for dynamic problems all without introducing non-physical artifacts into the solution, and its non-intrusive implementation into existing codes [23, 24].
Building on our past work in developing the Schwarz alternating method as a means to couple together FOMs [23, 24], intrusive projection-based ROMs [3] and physics-informed neural networks (PINNs) [38], we focus our attention herein on advancing the method to work with a non-intrusive model order reduction (MOR) technique known as operator inference (OpInf) [26, 36]. Unlike traditional intrusive MOR, which requires access to the underlying FOM code in order to project the governing PDE(s) onto a reduced subspace, OpInf works by assuming a functional form (usually linear or quadratic [10]) for the ROM in terms of to-be-learned reduced operators, and solving an optimization problem offline for these operators. This procedure significantly reduces both the development time and the time-to-impact. While DD-based couplings between ROMs and FOMs are not new, the majority of the literature on this topic has focused on developing intrusive coupling methods (e.g., Lagrange multipliers, optimization-based coupling) used to couple intrusive ROMs; the interested reader is referred to [3, 20, 5, 14, 7, 8, 27] and the references therein for more details. Related past work on data-driven couplings using Schwarz-like methods has focused on intrusive ROMs [6, 28, 17, 16], or on utilizing the coupling to accelerate NN training [38, 18, 19]. The proposed approach is most similar to the recent work by Farcas et al. [9], which develops a DD-based coupling of subdomain-local OpInf ROMs by learning appropriate reduced operators responsible for the coupling. Since each subdomain problem is solved once in [9], rather than by performing an iteration to convergence as done within our Schwarz framework, the subdomain-local solutions must be extended to the full domain and smoothly combined to achieve a continuous solution. Due to this, the resulting extended solutions can exhibit noticeable errors near the subdomain boundaries.
The remainder of this paper is organized as follows. In Section 2, we describe the overlapping version of the Schwarz alternating method applied to our model problem, the two-dimensional (2D) unsteady heat equation. In Section 2.2, we present some OpInf preliminaries. In Section 3, we describe our software implementations of the proposed Schwarz-based coupling approach applied to OpInf ROMs, hereafter called the OpInf-Schwarz method, which makes use of the open source FEniCSx [2] and OpInf [21] libraries. Results are presented in Section 4. We end with a concluding summary and a discussion of future work in Section 5.
2 The Schwarz Alternating Method Applied to Operator Inference ROMs
In the following sections, we will consider the specific model problem of the 2D heat equation, towards addressing the challenges that are encountered in this novel combination of the Schwarz method and operator inference. We stress that neither of these techniques inherently require an assumption of linearity, noting that the authors of the original operator inference paper [26] directly contrasted their method with the linearity assumption of Dynamic Mode Decomposition (DMD) [32]. While non-linearity will likely present further challenges to overcome for the OpInf-Schwarz coupling method, the heat equation provides an initial unsteady test problem that can suggest the feasibility of the method for more complicated problems.
2.1 Schwarz Alternating Method Preliminaries
Figure 1: Illustration showing an overlapping domain decomposition of a 2D domain $\Omega$ for the application of the Schwarz alternating method. Note $\Gamma_{1},\Gamma_{2}\not\subset\partial\Omega$.
Consider the heat equation specified as,
$\begin{array}[]{rcrl}\dot{u}(x,t)-\Delta u(x,t)&=&f(x),&\text{in }\Omega\times[0,T],\\ u(x,t)&=&g(x),&\text{on }\partial\Omega\times[0,T]\\ u(x,0)&=&v(x),&\text{in }\Omega,\end{array}$ (1)
where $\Omega\in\mathbb{R}^{p}$ is an open bounded domain for $p=1,2,3$ with boundary $\partial\Omega$, $f(x)$ and $g(x)$ are given source and boundary condition functions, respectively, $v(x)$ defines the initial condition, and $T>0$. Suppose we decompose the domain $\Omega$ into two overlapping subdomains $\Omega_{1}$ and $\Omega_{2}$, such that $\Omega_{1}\cup\Omega_{2}=\Omega$ and $\Omega_{1}\cap\Omega_{2}\neq\emptyset$ and, for later use, define $\overline{\Omega_{1}}$ as the closure of $\Omega_{1}$ and similarly for $\Omega_{2}$. Suppose also that we decompose the time interval $[0,T]$ into a set of non-overlapping time intervals $I_{n}=[t_{n},t_{n+1}]$, where $T\geq t_{n+1}>t_{n}\geq 0$. The overlapping Schwarz algorithm solves the following sequence of subdomain-local problems:
$\left\{\begin{array}[]{rcll}\dot{u}_{1}^{(k+1)}-\Delta u_{1}^{(k+1)}&=&f^{(k+1)},&\text{in }\Omega_{1}\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{1}})\times[t_{n},t_{n+1}]\\ u_{1}^{(k+1)}&=&u_{2}^{(k)},&\text{on }\Gamma_{1}\times[t_{n},t_{n+1}],\end{array}\right.$ (2)
and
$\left\{\begin{array}[]{rcll}\dot{u}_{2}^{(k+1)}-\Delta u_{2}^{(k+1)}&=&f^{(k+1)},&\text{in }\Omega_{2}\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&g,&\text{on }(\partial\Omega\cap\overline{\Omega_{2}})\times[t_{n},t_{n+1}]\\ u_{2}^{(k+1)}&=&u_{1}^{(k+1)},&\text{on }\Gamma_{2}\times[t_{n},t_{n+1}],\end{array}\right.$ (3)
for Schwarz iteration $k=0,1,2,...$ subject to initial conditions $u_{1}(x,0)=v|_{\Omega_{1}}$ and $u_{2}(x,0)=v|_{\Omega_{2}}$. In (2), $u_{i}$ for $i=1,2$ denotes the solution in subdomain $\Omega_{i}$, and $\Gamma_{i}$ is the so-called Schwarz boundary (see Figure 1). It is common to set $u_{1}^{(0)}=v|_{\partial\Omega_{1}}$ and $u_{2}^{(0)}=v|_{\partial\Omega_{2}}$ when initializing the Schwarz iteration process, to ensure solution compatibility with the initial condition. A key advantage of the Schwarz alternating method is that it allows for the subdomains $\Omega_{1}$ and $\Omega_{2}$ to be discretized using different meshes and/or element types [23, 24]. In the case where $\Omega_{1}$ and $\Omega_{2}$ are discretized using different meshes and do not have a coincident interface, applying the Schwarz boundary condition on the Schwarz boundaries $\Gamma_{i}$ requires the construction of a projection operator; this can be done via a simple application of finite element interpolation functions readily available in most codes [24, 22]. The iterative process in (2) and (3) continues until a set of pre-determined criteria are met. In the present work, convergence criteria are based on the Euclidian norm of the solution differences between consecutive Schwarz iterations; that is, Schwarz is deemed converged when $\epsilon_{\text{abs}}^{(k)}<\delta_{\text{abs}}$ and $\epsilon_{\text{rel}}^{(k)}<\delta_{\text{rel}}$ for some pre-specified Schwarz tolerances $\delta_{\text{abs}},\delta_{\text{rel}}>0$, where
$\epsilon_{\text{abs}}^{(k)}:=\sqrt{||u_{1}^{(k)}-u_{1}^{(k-1)}||^{2}+||u_{2}^{(k)}-u_{2}^{(k-1)}||^{2}},$ (4)
and
$\epsilon_{\text{rel}}^{(k)}:=\sqrt{\frac{||u_{1}^{(k)}-u_{1}^{(k-1)}||^{2}}{||u_{1}^{(k)}||^{2}}+\frac{||u_{2}^{(k)}-u_{2}^{(k-1)}||^{2}}{||u_{2}^{(k)}||^{2}}},$ (5)
for Schwarz iteration $k=1,2,...$. The Schwarz iteration process (2)–(3) is converged within each time interval $[t_{n},t_{n+1}]$ before moving on to the next time interval. A key advantage of this time-stepping approach is that different time integrators and time steps can be used in different subdomains; for a detailed discussion of Schwarz time-stepping and related machinery, the reader is referred to [24, 22]. In the numerical experiments presented herein, we restrict attention to the case where all subdomains have the same time integrator and time step, as our goal is to assess the method’s viability when coupling subdomain-local operator inference-based ROMs with each other and with subdomain-local FOMs while eliminating other confounding factors.
2.2 Operator Inference Preliminaries
Operator inference is a data-driven, non-intrusive, projection-based method for the model order reduction of PDEs, developed relatively recently by Peherstorfer and Willcox [26] as an alternative to standard Galerkin projection ROMs. It consists of two major steps:
  1. 1.
    Build a reduced basis from data.
  2. 2.
    Infer operators on the span of this reduced basis.
These two steps are described succinctly below.
2.2.1 Proper Orthogonal Decomposition
To construct a reduced basis, we will use proper orthogonal decomposition (POD) [13, 37], also known as principal component analysis (PCA) in some fields, though one might also use other methods, e.g., the reduced basis method with a greedy algorithm [30]. To perform POD, we require snapshots from some data source or FOM. Continuing the example of the heat equation from (1), a spatially discretized FOM for the heat equation will typically appear in the following form after a boundary lift:
$\dot{\mathbf{x}}=\mathbf{K}\mathbf{x}+\mathbf{B}\mathbf{g}+\mathbf{f},$ (6)
where $\mathbf{x}\in\mathbb{R}^{N}$ and $\mathbf{f}\in\mathbb{R}^{N}$ are now discretized vectors corresponding to the unconstrained state degrees of freedom (DoFs), and $\mathbf{g}\in\mathbb{R}^{m}$ discretizes the Dirichlet boundary condition. The matrix $\mathbf{K}\in\mathbb{R}^{N\times N}$ comes from discretizing the continuous Laplace operator $\Delta$, and $\mathbf{B}\in\mathbb{R}^{N\times m}$ deals with effects of the boundary term. The full state representation $\mathbf{u}\in\mathbb{R}^{N+m}$ for all DoFs is obtained by augmenting the unconstrained solution $\mathbf{x}$ with the known boundary condition $\mathbf{g}$.
After (6) has been solved in time for $\tau$ separate states inclusively between time $t=0$ and $t=T$ (i.e., $0=t_{1}<t_{2}<\dots<t_{\tau}=T$), we obtain a collection of unconstrained state snapshots,
$\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{\tau}]\in\mathbb{R}^{N\times\tau},$ (7)
[…]
$\max_{\tilde{\psi}\in H}\sum_{i=1}^{\tau}\frac{\left|\langle\mathbf{x}_{i},\tilde{\psi}\rangle\right|^{2}}{\|\tilde{\psi}\|^{2}}.$ (8)
Equation (8) is satisfied by $\psi_{1}$, the left singular vector associated with $\sigma_{1}$, the largest singular value of $\mathbf{X}$ as shown in [40]. Further basis vectors up to dimension $r\leq d$ are chosen to be orthonormal to previous basis vectors, i.e., $\langle\mathbf{\psi}_{i},\psi_{j}\rangle=\delta_{ij}$, the Kroneker delta. The columns of $\mathbf{\Psi}$ necessarily satisfy this additional property. The restriction of $\mathbf{\Psi}$ to its first $r$ columns will be referred to as $\mathbf{\Psi}_{r}$ going forwards.
[…]
2.2.2 Operator Inference
Operator inference is a form of projection-based model that uses a POD basis to reduce the dimension of the operators assembled for the FOM, which, in this paper, is the heat equation described in (6). In a classical Galerkin ROM, the operator $\mathbf{K}_{r}\in\mathbb{R}^{r\times r}$ is the operator which best represents the action of $\mathbf{K}$ in the reduced space of dimension $r$ defined by the basis $\mathbf{\Psi}_{r}$. That is,
$\mathbf{K}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{K}\mathbf{\Psi}_{r}.$
This is intrusive because it requires access to the FOM matrix $\mathbf{K}$, and, practically, the code that produced $\mathbf{K}$. Similarly, we can obtain $\mathbf{x}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{x}$ and $\mathbf{f}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{f}$, both belonging to $\mathbb{R}^{r}$.
The key difference between operator inference and traditional intrusive projection-based MOR is that, in the former approach, the reduced operators are not created via intrusive projection, but inferred directly using available snapshot data. OpInf is based on the observation that a projection-based ROM derived from a FOM with polynomial nonlinearities will possess the same algebraic structure as the FOM.
For the problem considered herein, the 2D heat equation, it is straightforward to see that a projection-based ROM will preserve the linear algebraic structure of the FOM, (6). In order to obtain the reduced operators in (10), one solves a regression problem for these reduced operators, $\hat{\mathbf{K}}\in\mathbb{R}^{r\times r}$ and $\hat{\mathbf{B}}\in\mathbb{R}^{r\times m}$. In particular, given $j<\tau$ steps of FOM training data $\mathbf{x}_{i}$, $i=1,\dots,j$, minimize,
$\min_{\hat{\mathbf{K}},\hat{\mathbf{B}}}\sum_{i=1}^{j}\|\dot{\mathbf{x}}_{r,i}-\hat{\mathbf{K}}\mathbf{x}_{r,i}-\hat{\mathbf{B}}\mathbf{g}_{i}-\mathbf{f}_{r,i}\|_{2}^{2}.$ (9)
The subscript $r$ is here to note the components in the ROM dimension which are explicitly known within the training data. In the case that a technique known as re-projection is used, it can be shown that the learned OpInf ROM system recovers the standard Galerkin projected intrusive ROM [25]. If $\dot{\mathbf{x}}_{r,i}$ is not explicitly available, it may be estimated from available data using a difference method. The hat notation is reserved for those components which belong to the OpInf method. As this is a data-driven minimization problem, it is not strictly necessary to use the actual known boundary condition $\mathbf{g}_{i}$ or forcing function $\mathbf{f}_{r,i}$; some approximation or ansatz can be learned instead, and/or these terms can be combined into a single input term.
Once problem (9) is solved, which in implementation will typically involve a QR based approach. The ROM ODE analogue of (6) takes the form
$\dot{\hat{\mathbf{x}}}=\hat{\mathbf{K}}\hat{\mathbf{x}}+\hat{\mathbf{B}}\mathbf{g}+\mathbf{f}_{r},$ (10)
where we now solve for the reduced state variable $\hat{\mathbf{x}}\in\mathbb{R}^{r}$ because the outputs of this ODE are not guaranteed to be the same as those for the standard intrusive Galerkin projection. It is guaranteed that, under the assumptions that the time-stepping scheme is convergent as $dt\to 0$, the approximation $\dot{\hat{\mathbf{x}}}_{j}$ converges to the true time derivative at time $t_{j}$ as $dt\to 0$, and linear independence of supplied data ($d\geq\tau$, the total number of snapshots), then, for all $0<\epsilon\in\mathbb{R}$ there exists a timestep $dt$ for the FOM and a $r\leq d$ such that,
$\|\mathbf{K}_{r}-\hat{\mathbf{K}}\|_{F}<\epsilon.$
The same result holds for $\hat{\mathbf{B}}$ [26].
While this overview has focused on the heat equation, we again stress that operator inference is general to low-order polynomial non-linearity, with results previously shown for a polynomial non-linearity of degree 3 in a one-dimensional (1D) nuclear reactor model in [26], for a 2D single injector combustion model in [36], and for a three-dimensional (3D) rotating detonation rocket engine in [9].
3 Software Implementations
The Schwarz coupling implementations we will test are FOM-ROM and ROM-ROM, where FOM refers to a finite element (FE) simulation and ROM to operator inference. We will draw our data source from a monolithic FE simulation, but the FOM-ROM coupling will have its simulation run from scratch on its own subdomain independently of the monolithic simulation.
We use the open source FE library FEniCSx in Python for all FE simulations. FEniCSx is a FE problem solving environment with the ability to define variational forms close to actual mathematical notation [1], along with various tools [2, 35, 34] to enable the creation of a FE space on which computations can be performed. Boundary conditions are enforced strongly through a lifting method leading to a FOM system similar to (6).
Operator inference is performed with the Python library OpInf, implemented by Shane McQuarrie [21]. OpInf is a set of tools to carry out POD and infer the required operators as described in Section 2.2. The required data will be drawn from a monolithic FE simulation on $\Omega$, also performed in FEniCSx. When solving the OpInf regression problem (9), we estimate the required state derivative information $\dot{\hat{\mathbf{x}}}$ for through a first order backward difference of available state data, and our ROM ODE (10) will be solved using backward Euler.
The numerical stability of OpInf models is a significant concern and area of active research. Operator inference models are not guaranteed to be structure-preserving [11], and one key structure that we care about is the theoretical stability of the ODE system (10). FOM solutions are generated presently through the FE method, and as such the eigenvalues of $\mathbf{K}$ in (6) are guaranteed to be negative. This structure is not guaranteed to be passed on to $\hat{\mathbf{K}}$, and often is not. There are a variety of strategies that can help improve stability of reduced models – independently of OpInf, data pre-processing is commonly used in ROMs to create a model for a transformed set of snapshots. We will typically use a centering approach where we build a model for snapshot data $\mathbf{x}-\mathbf{\bar{x}}$, the mean-centered snapshots.
Specific to OpInf, we also employ a regression regularization strategy for problem (9). In literature [26], the basic OpInf problem is typically cast into the form,
$\min_{\mathbf{O}}\|\mathbf{D}\mathbf{O}^{T}-\mathbf{R}^{T}\|_{F},$ (11)
[…]
Finally, the Schwarz method that couples these problems has been implemented by the authors of this paper in accordance with Section 2.1. To simplify the boundary transmission of the Schwarz process, only conformal meshes will be used for the differing domains. We have also chosen for this first test to implement the overlapping Schwarz method with a constant time step $dt$ in all domains. Neither of these two specifications are required constraints on the Schwarz method.
4 Results
[…]
For the purposes of an initial test of the OpInf-Schwarz method, we are most interested if the model matches the standard behavior of ROMs and the Schwarz method. We expect that, as the ROM dimension $r$ and the data allowance increases, the error should approach 0, and that as we increase the size of the Schwarz overlap, the number of Schwarz iterations to converge per timestep will decrease. Lastly, as for any ROM, we expect a significant speed-up in computation time for the price we pay in accuracy.
$\Omega_{1}$$\Omega_{2}$
(a) Vertical configuration
$\Omega_{1}$$\Omega_{2}$
(b) Horizontal configuration
$\Omega_{1}$$\Omega_{2}$$\Omega_{3}$$\Omega_{4}$
(c) 4 Squares configuration
Figure 2: Overlapping domain decomposition configurations
We will investigate each of these points in a variety of domain decomposition scenarios according to Figure 2. For a global domain, we choose the square $\Omega=[-1,1]\times[-1,1]\in\mathbb{R}^{2}$ on which a monolithic FE simulation will be carried out. The monolithic boundaries will be referred to as $\partial\Omega_{L}$ for the left, $\partial\Omega_{R}$ for the right, $\partial\Omega_{T}$ and $\partial\Omega_{B}$ for the top and bottom of the square respectively. We will divide each square into either two rectangles or four squares, each with an equal Schwarz overlap measured in the number of elements overlapping each neighboring subdomain. For these Schwarz overlaps, explicit communication is only between neighbors which share at least one complete edge – for example, boundary information can be communicated to a left or right neighbor, but not a diagonal neighbor. The Schwarz subdomains will be updated in ascending numerical order.
Aside from the consistent monolithic domain, we will also use the following consistent choices across all tests. We will be using the heat equation as specified in (1) with $f=0$. Our time domain will be $[0,1]$, our time step will be $dt=0.01$ leading to 101 steps, inclusively, from the initial condition at $t=0$ to the final evaluation at $t=1$. The monolithic domain will be split into a 50 by 50 grid, and then triangular Lagrangian elements of degree 1 will be formed by splitting the grid diagonally. These choices lead to 2601 FE nodes on the monolithic mesh, inclusive of the boundary. Within the Schwarz method, our relative (5) and absolute (4) tolerances for convergence will both be fixed at $1\times 10^{-10}$.
Errors
Our typical measure of error for OpInf-Schwarz is the average pointwise relative error, i.e.,
$\mathbf{E}_{\ell^{2}}^{\text{avg}}=\frac{1}{\tau-1}\sum_{i=2}^{\tau}\frac{\|\mathbf{u}_{i,\text{MONO}}-\mathbf{u}_{i,\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}_{i,\text{MONO}}\|_{2}},$
Where $\mathbf{u}_{i,\text{MONO}}$ refers to the finite element monolithic simulation carried out over $\Omega$ at time step $i$ and $\mathbf{u}_{i,\text{SCHWARZ}}$ to the solution via the Schwarz method merged from the component subdomains at time step $i$. The norm $\|\cdot\|_{2}$ refers to the $\ell^{2}$ norm, and both $\mathbf{u}_{i,\text{MONO}}$ and $\mathbf{u}_{i,\text{SCHWARZ}}$ are vectors of solution values at each FE node. The error determination is made without including the initial condition, as they are set to be equal.
The projection error of the snapshots onto the POD basis in subdomain $\Omega_{j}$ (dependent on basis dimension $r$) is here defined as,
$\mathbf{E}_{\text{proj}}^{\Omega_{j}}=\frac{\|\mathbf{X}_{j}-\mathbf{\Psi}_{r,j}\mathbf{\Psi}_{r,j}^{T}\mathbf{X}_{j}\|_{F}}{\|\mathbf{X}_{j}\|_{F}},$
where $\mathbf{X}_{j}$ are the snapshots associated with all solved for DoFs in $\Omega_{j}$ over the entire time domain, and $\mathbf{\Psi}_{r,j}$ is the $r$-dimensional basis associated with the solved for DoFs of $\Omega_{j}$. This gives a measure of the approximation quality of the POD basis in each subdomain. When averaged across all subdomains, this will be reported using the symbol, $\mathbf{E}_{\text{proj}}^{\text{avg}}$.
Lastly, we will also use the maximum error in 4.2, defined as,
$\mathbf{E}_{\ell^{2}}^{\text{max}}=\max_{i}\frac{\|\mathbf{u}_{i,\text{MONO}}-\mathbf{u}_{i,\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}_{i,\text{MONO}}\|_{2}}.$
Quantities Reported in Tables
In addition to errors, we will be reporting the following quantities in following tables.
  1. 1.
    “Avg S.I.” is the mean number of Schwarz iterations required to reach convergence across all time steps.
  2. 2.
    “Overlap” is the amount by which the $50\times 50$ grid is shared by neighboring subdomains before being broken up into triangular elements.
  3. 3.
    “Time” is the average wall clock time taken to run the online model on identical hardware, measured in seconds. The number of runs averaged will be mentioned in each table. This does not include any time spent setting up matrices, generating data, or similar offline tasks.
  4. 4.
    $r$” is the integer dimension of the OpInf basis generated through POD. This is uniform across multiple subdomains.
  5. 5.
    “Data” is the number of training steps taken, with a value of 1 only including the initial condition. There are a maximum of 101 steps possible across the time interval $[0,1]$, and any value of “Data” less than 101 will result in prediction outside of training data. “Data” is uniform across all subdomains.
Lastly, computational speed-up will be measured as the wall clock time to run the online portion of the OpInf-Schwarz coupled ROM models relative to that of the fully FE coupled Schwarz method and the monolithic FE solution over the same domain configuration and on the same hardware.
4.1 Static Boundary Conditions
For a first test, we consider the case where $u(\partial\Omega_{T},t)=u(\partial\Omega_{B},t)=0$, $u(\partial\Omega_{L},t)=2$ and $u(\partial\Omega_{R},t)=5\,\,\forall\,t\in(0,1]$. The discontinuity at some of the corners is resolved in favor of the top and bottom boundaries. We establish a baseline in Table 1 by coupling solely FE models in each of the 3 configurations of Figure 2. The wall time, in seconds, was averaged over 5 separate runs. Each subdomain has an overlap of 10 with its non-diagonal neighbors.
$\mathbf{E}_{\ell^{2}}^{\text{avg}}$ $1.27\times 10^{-14}$
Avg S.I. 4.41
Time (s) 3.03
(a) Vertical, Fig 2(a)
$\mathbf{E}_{\ell^{2}}^{\text{avg}}$ $1.57\times 10^{-14}$
Avg S.I. 4.32
Time (s) 3.59
(b) Horizontal, Fig 2(b)
$\mathbf{E}_{\ell^{2}}^{\text{avg}}$ $1.94\times 10^{-14}$
Avg S.I. 5.9
Time (s) 7.05
(c) 4 Squares, Fig 2(c)
Table 1: Baseline for all-FE coupled models, Overlap = 10
The full monolithic simulation, averaged over 5 runs, took 0.94 seconds to run the online portion, and by definition has $\mathbf{E}_{\ell^{2}}^{\text{avg}}=0$.
Vertical Configuration
In Figure 3 and Table 2, we examine the behavior of the OpInf-Schwarz method with respect to $r$, the subdomain overlap, and data by coupling two OpInf models in the vertical orientation of Figure 2(a). Wall clock time results, in seconds, are averaged over 2 individual runs. We will additionally be using a centering strategy on the snapshots, but not scaling the snapshots in any way.
(a) 2 coupled OpInf models in vertical configuration, $t=0$
(b) Error between OpInf-Schwarz and monolithic FE simulation over time.
Figure 3: OpInf-Schwarz vs. monolithic simulation. OpInf parameters: $r=6$, Overlap = 10, 30 steps training data.
Data $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ Avg S.I. Time
10 $1.16\times 10^{-01}$ 4.0 0.30
20 $5.41\times 10^{-03}$ 4.57 0.30
30 $6.02\times 10^{-04}$ 4.11 0.27
50 $1.56\times 10^{-04}$ 4.52 0.31
80 $9.16\times 10^{-05}$ 4.39 0.28
100 $8.62\times 10^{-05}$ 4.49 0.31
(a) Fixed parameters: Overlap = 10, $r=6$.
Overlap $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ Avg S.I. Time
1 $8.68\times 10^{-03}$ 6.4 0.35
2 $5.33\times 10^{-03}$ 6.34 0.34
5 $2.36\times 10^{-03}$ 5.7 0.31
10 $6.02\times 10^{-04}$ 4.11 0.22
20 $5.96\times 10^{-04}$ 4.7 0.25
40 $8.18\times 10^{-04}$ 4.26 0.24
(b) Fixed parameters: 30 steps of training data, $r=6$.
$r$ $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ Avg S.I. $\mathbf{E}_{\text{proj}}^{\text{avg}}$ Time
2 $2.34\times 10^{-02}$ 8.04 $8.98\times 10^{-02}$ 0.31
4 $4.24\times 10^{-03}$ 5.0 $1.02\times 10^{-02}$ 0.25
6 $6.02\times 10^{-04}$ 4.11 $6.14\times 10^{-04}$ 0.19
8 $1.74\times 10^{-03}$ 5.0 $5.85\times 10^{-05}$ 0.24
10 $1.85\times 10^{-03}$ 5.0 $6.02\times 10^{-06}$ 0.23
30 $1.85\times 10^{-03}$ 5.0 $3.54\times 10^{-10}$ 0.25
(c) Fixed parameters: 30 steps of training data, Overlap = 10.
Table 2: Comparison of OpInf-OpInf coupled models in vertical orientation (Fig 2(a)).
From a ROM perspective, Table 2(c) has several interesting results which deserve explanation. First, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ only decreases up to $r=6$, then increases before stalling out for higher values of $r$. This is despite the fact that the projection error of the snapshots onto the basis drops consistently with the basis size. In the row $r=6$, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ and $\mathbf{E}_{\text{proj}}^{\text{avg}}$ have the same order of accuracy, while $\mathbf{E}_{\text{proj}}^{\text{avg}}$ is lower for all further values of $r$. This suggests that the accuracy of the OpInf-Schwarz method is limited by the basis quality up until about $r=6$, and afterwards is most limited by optimization error in the inference step. A basis size of $r=3$ and $r=4$ is necessary to capture 99.9% of the total energy of the snapshots $\Omega_{1}$ and $\Omega_{2}$ respectively.
The second item of note in Table 2(c) is that the average run time goes down by about a third as we increase $r$ from 2 to 6, even though we are increasing the size of the linear systems we are solving. This is due to the coupling of poor quality solutions when $r=2$, which requires additional Schwarz iterations to converge as shown by the mean Schwarz iterations. A similar phenomenon is observed in [3] when coupling subdomain-local intrusive ROMs via the Schwarz alternating method.
The remaining subtables of Table 2 reveal that OpInf-Schwarz is working in the expected manner. In Table 2(b), increasing the size of the overlap decreases the number of Schwarz iterations, and in Table 2(a), increasing the training data decreases the approximation error. Since we are using the same data for both the POD basis and the OpInf training, it is fair to ask whether increasing the data allowance also allows POD to extract useful data for larger values of $r$ – that is, further decreases to $\mathbf{E}_{\text{proj}}^{\text{avg}}$ In our testing, increasing training data had a minor effect on $\mathbf{E}_{\text{proj}}^{\text{avg}}$ compared to increasing $r$, though this would likely become more relevant in more challenging problems. The minimal error we were able to recover was $\mathbf{E}_{\ell^{2}}^{\text{avg}}=2.5\times 10^{-05}$ with the choices of $r=10$, Overlap = 10, and 100 steps of training data.
While the times recorded are certainly not deterministic, a comparison between the FE-FE coupling of Table 1(a) and the minimum and maximum values of Table 2 suggest a speed-up of 8.6 to 15.2 times for fully OpInf-OpInf coupled models as compared to a FOM-FOM coupling on the same spatio-temporal grid, and 2.7 to 4.9 times speed-up for fully OpInf-OpInf coupled models as compared to the monolithic FE simulation. Further speed-ups are possible by introducing the “additive” Schwarz method [41], which computes subdomain solutions in parallel.
Comparison With Other Configurations
We now investigate whether interactions between the subdomain geometry and boundary conditions have an effect on the performance of OpInf-Schwarz. Because the non-zero boundary conditions in this simple case are on the left and right, a vertical overlap splitting the middle has a less complicated gradient than a horizontal overlap. In the more realistic scenario of $r=6$, Overlap = 10, and 30 steps of training data, Table 3 shows that the error and Schwarz iterations are mildly worse for those configurations that impose a more complicated Schwarz boundary.
Model Configuration $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ Avg S.I. Time (s)
Vertical 6.02 ×$\times 10^{-04}$ 4.11 0.26
Horizontal 8.91 ×$\times 10^{-04}$ 4.17 0.26
4 Squares 7.10 ×$\times 10^{-04}$ 5.81 0.61
Table 3: Vertical, horizontal, and square configurations for simple boundaries
All of the results in the simple case of Section 4.1 are in line with prior expectations for Schwarz performance in a ROM setting.
4.2 Time-Varying Boundary Conditions
We now consider a more challenging scenario for the 2D heat equation. Instead of solely constant boundary conditions, define
$v(t,\mu)=1+0.5\sin(2\pi\mu t),$ (12)
[…]
The offset frequencies of the boundary conditions on the left and top boundaries create dynamics which evolve through a significant portion of the time domain, especially in the top left quadrant. These types of problems can be challenging for ROMs to resolve when given insufficient data, though they can usually be fixed by increasing the training data. This is one of the essential choices in ROMs – increasing the training range can often greatly improve accuracy, but data generation in realistic problems can be enormously expensive, especially if it must be generated from a monolithic simulation. Domain decomposition frameworks provide options beyond increasing the training range, and we will now consider a comparison between purely OpInf-Schwarz with a FE-OpInf-Schwarz coupled problem.
4 Squares Configuration
Figure 4 shows two OpInf-Schwarz solutions at the final time step of $t=1$ in the four square configuration of Figure 2(c). For the OpInf portions, each solution has the same ROM basis dimension $r=6$, 40 steps of training data (out of a possible 101), and a mean centering strategy performed on the snapshots. They additionally share the same boundary and initial conditions as described above. The only difference in their setup is that Figure 4(a) couples four OpInf models, while Figure 4(b) has replaced one of the OpInf models with a FE model in $\Omega_{3}$, the top-left subdomain.
The FE assimilation in Figure 4(b) is not perfect if one looks closely along $x=0$ or $y=0$, but it shows a cohesive solution very similar to the monolithic solution in 4(d), while the solution in Figure 4(a) is clearly failing and is in fact about to diverge entirely. The fully OpInf model with a higher data allowance in Figure 4(c) is comparable to Figure 4(b), and both fail to resolve some aspects of the monolithic simulation.
(a) OpInf solution in all 4 subdomains, 40 steps of training data
(b) FE solution in top left $\Omega_{3}$ only, 40 steps of training data for OpInf portions
(c) OpInf solution in all 4 subdomains, 65 steps of training data
(d) Monolithic FOM (Finite Element Simulation)
Figure 4: OpInf-Schwarz solutions and Monolithic FOM at $t=1$.
One interesting feature of Figure 4(a) is that the Schwarz communication pattern is clearly evident. The solution is clearly incorrect in the top left subdomain, $\Omega_{3}$. The error in $\Omega_{3}$ subsequently propagates to neighboring $\Omega_{1}$ and $\Omega_{4}$, while the bottom right subdomain $\Omega_{2}$ remains closest to an accurate solution.
Following from the results in Table 4, a purely OpInf coupled Schwarz model would require 65 total timesteps (an additional 25 steps of training data) to exceed the average relative accuracy of the OpInf-Schwarz with one FE subdomain and 40 steps of training data (as is displayed in Figure 4(b)).
Data $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ $\mathbf{E}_{\ell^{2}}^{\text{max}}$
30 $1.97\times 10^{-01}$ $1.22\times 10^{00}$
35 $5.41\times 10^{-01}$ 4.07×$\times 10^{-01}$
40 $1.07\times 10^{-01}$ $6.40\times 10^{-01}$
50 $3.62\times 10^{-02}$ $2.12\times 10^{-01}$
60 $9.15\times 10^{-03}$ $3.13\times 10^{-02}$
65 $6.48\times 10^{-03}$ $1.91\times 10^{-02}$
(a) All OpInf, 4 Coupled OpInf Models
Data $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ $\mathbf{E}_{\ell^{2}}^{\text{max}}$
30 $3.33\times 10^{-02}$ $7.06\times 10^{-02}$
35 $9.47\times 10^{-03}$ $2.75\times 10^{-02}$
40 $7.14\times 10^{-03}$ $1.89\times 10^{-02}$
50 $4.76\times 10^{-03}$ $1.11\times 10^{-02}$
60 $1.86\times 10^{-03}$ $1.4\times 10^{-02}$
65 $1.77\times 10^{-03}$ $1.40\times 10^{-02}$
(b) 3 OpInf 1 FE, FE model in top left subdomain $\Omega_{3}$ only.
Table 4: OpInf-Schwarz errors for various training data ranges. $r=6,$ Overlap = 10.
Of course, this stability comes at a price: over two separate runs, the average time taken for the 4 coupled OpInf models in Table 4(a) was 0.73 seconds, while the average for the substitution of one FE model in $\Omega_{3}$ in Table 4(b) is 3.15 seconds, a 4.3 times slowdown over a purely OpInf coupled model. The Schwarz framework gives us the opportunity to more evenly weigh the opportunity cost for amount of data required vs. computational speed for specific problems.
In comparison, a completely FE coupled model with the same settings has $\mathbf{E}_{\ell^{2}}^{\text{avg}}=\mathcal{O}(10^{-14})$, but, over three runs, took an average of 7.9 seconds. A monolithic FE sim takes approximately 0.8 seconds. The “All OpInf” model is therefore about 9 times faster, and the 3 OpInf 1 FE model about 2.4 times faster than the fully FE coupled model. Comparing against the monolithic FOM, the “All OpInf” model is 1.1 times faster and the 3 OpInf 1 FE is 3.9 times slower.
Efficiency Comparison With Other Configurations
We summarize our timing results for the time-varying boundary problem in the Pareto plot of Figure 5 using data from Table 4 for the “4 Squares” labelled results, and additional results for the “Vertical” labelled results produced in the vertical configuration from Figure 2(a). In particular, “Vertical FE-OpInf” was produced with the FE model in the left subdomain, and “Horizontal FE-OpInf” was produced with the FE subdomain as the top subdomain, both advantageous DDs for resolving the more complicated dynamics in the top left portion of $\Omega$. Fixed parameters are $r=6$ and Overlap = 10. This chart demonstrates the significant impact that subdomain geometry can have when interacting with more complicated boundary issues or solution features. The vertical OpInf-OpInf is more accurate for smaller data allowances, but it is overtaken in accuracy by the 3 OpInf 1 FE model for higher data allowances. All of the models with a FE component have similar time costs, but the horizontal and 4 square configurations make better use of additional data than the vertical configuration does.
Figure 5: Pareto plot for time-varying BCs, $\mathbf{E}_{\ell^{2}}^{\text{avg}}$ vs Time (s). Fixed parameters: $r=6$, Overlap $=10$. Plotted points are “Data” as it ranges over 30, 40, 50, and 60 as in Table 4. Values in bottom left are preferred.
While none of the models incorporating a FE subdomain exhibit lower runtimes than that of the monolithic FOM, we reiterate that it is possible to improve the performance of the Schwarz iteration process via the additive Schwarz formulation.
5 Conclusions and Future Work
A major impetus driving the development of the OpInf-Schwarz method is the desire for a minimally intrusive methodology for constructing ROMs that can interface modularly with existing high performance codes. A truly data-driven and modular domain decomposition-based ROM has the potential to greatly reduce computational expenses incurred by meshing and re-meshing complicated 3D objects, and therefore in performing long run-time and/or multi-query simulations.
The preliminary investigation presented herein has assessed various characteristics of the OpInf-Schwarz framework in the context of a 2D heat equation. Our results have shown that the OpInf-Schwarz method is capable of coupling between both OpInf and FE subdomains to recover accurate solutions, and that a purely OpInf formulation can run faster than a monolithic simulation. We have also demonstrated that OpInf-Schwarz is very flexible, and can be tuned in several ways to the problem at hand.
At the same time, this work has uncovered several challenges that are worth investigating in the future. First, the calculated average relative pointwise error presented in Section 4.1 stagnates at around $\mathcal{O}(10^{-5})$ and does not decrease further despite a increase to the POD basis size and training data. While this level of error is acceptable for many applications, it would be ideal for the method to be truly convergent to the data source. While it seems likely that the issue lies with the optimization error of the regression in problem (9) given that the projection errors reported in Table 2(c) continue to decrease, a comparison with a standard intrusive projection ROM would confirm whether the cause is optimization error or errors induced by the coupling process. It may be possible to reduce optimization errors by employing a technique called re-projection [25], and/or by improving the variational consistency of the model [11].
There are also potential improvements to the implementation of OpInf presented in this paper, as the stability of the learned operators is not enforced at the inference step. At present, this is resolved with a simple Tikhonov-based regularization strategy, but a more rigorous approach may be needed in more complicated problems, e.g., by using a more sophisticated regularization strategy such as those presented in [36, 31]. There are additional opportunities to improve the implementation of the Schwarz boundary conditions within our OpInf-Schwarz formulation in a way that both reduces computational cost and improves accuracy.
Currently, the learned boundary operator $\hat{\mathbf{B}}$ scales with the size of the boundary in our formulation but only transmits boundary information weakly at a ROM level into the interior of each subdomain. It is likely that a more sophisticated boundary enforcement strategy is needed. One promising approach is based on centering the reduced POD basis [11] (similar to what is commonly done for conventional intrusive ROMs [12]), as opposed to centering the snapshots, as was done in this paper. Reducing the computational complexity of our method requires reducing the number of columns of $\hat{\mathbf{B}}$, which can be large for multi-dimensional problems with many boundary nodes. To mitigate the cost associated with evaluating $\hat{\mathbf{B}}\mathbf{g}$, it may be possible to perform a separate dimension reduction of this term. Finally, as noted earlier, the overall computational complexity of OpInf-Schwarz can be improved by introducing “additive” Schwarz, characterized by parallel subdomain solves with asynchronous boundary condition communication [41].
The last major point is on extending the Opinf-Schwarz method to more challenging and realistic problems, as we expect the greatest utility of this model to be shown in more complicated situations. The heat equation can be revealing for a new model, but extending the model to 3D, non-linear, convection dominated, and parametric problems [39] is a priority, as are implementing strategies to choose optimal domain decomposition and model assignment to enable on-the-fly ROM-FOM switching for maximal performance.
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Mathematical Multifaceted Integrated Capability Centers (MMICCs) program, under Field Work Proposal 22-025291 (Multifaceted Mathematics for Predictive Digital Twins (M2dt)), Field Work Proposal 20-020467, and the Laboratory Directed Research and Development program at Sandia National Laboratories. The writing of this manuscript was funded in part by the fourth author’s (Irina Tezaur’s) Presidential Early Career Award for Scientists and Engineers (PECASE). This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan.
The authors would like to thank Dr. Shane McQuarrie for numerous helpful technical discussions in the area of operator inference, and for providing support on the OpInf library, used to generate the results presented in this paper.